OpenStudy (anonymous):

determine the coordinates of the midpoint of the line segment whose endpoints are A(0, 2, 4) and B(-8, 2, -6)

OpenStudy (lgbasallote):

wouldnt it just be \[\large (\frac{x_1 + x_2}{2} \; , \; \frac{y_1 + y_2}{2} \; , \; \frac{z_1 + z_2}{2})\]